OpenStudy (raffle_snaffle):

A particular telephone number is used to receive both voice calls and fax messages. Suppose that 25% of the incoming calls involve fax messages, and consider a sample of 25 incoming calls. (Round your answers to three decimal places.) (a) What is the probability that at most 4 of the calls involve a fax message? (b) What is the probability that exactly 4 of the calls involve a fax message? (c) What is the probability that at least 4 of the calls involve a fax message? (d) What is the probability that more than 4 of the calls involve a fax message?

OpenStudy (raffle_snaffle):

@mathmale

OpenStudy (raffle_snaffle):

I was up until midnight trying to figure out c and d. I got answers, but I am not understanding something I will show you when you get here

OpenStudy (raffle_snaffle):

okay so for part a...

OpenStudy (mathmale):

Hello! Am I correct in assuming that this is binomial probability? If so, what is n? p?

OpenStudy (raffle_snaffle):

give me a sec. I know I am suppose to use the binomial distribution but I don't want to worry about that now. Let me upload a pic

OpenStudy (raffle_snaffle):

So does P(X=4) = P(X<=4) - P(<=3) is this a true statement?

OpenStudy (mathmale):

I agree with your approach to Problem 11a. Before I answer your question: do you have a TI-83 or -84 calculator on hand?

OpenStudy (raffle_snaffle):

I have a TI-89 on hand

OpenStudy (raffle_snaffle):

I don't have an 83 or 84 just an 89.

OpenStudy (raffle_snaffle):

I got that answer from previous work. You want to see?

OpenStudy (mathmale):

Fancy, fancy. In my old (retired) age I have only a TI-83 to my name. But your TI-89 will surely calculate binomial probabilities. May i assume you've used it for that purpose before? No, your Problem 11a answer is clear, correct and appropriate.

OpenStudy (mathmale):

calculator will do both the binomial cumulative probability and the probability of ONE particular x value.

OpenStudy (raffle_snaffle):

Yes, I was shown from another tutor how to use nCy(y,n) command.

OpenStudy (raffle_snaffle):

when you say cumulative you mean continuous?

OpenStudy (mathmale):

my calc has these 2 commands that are relevant here: binompdf and binomcdf. You familiar with both?

OpenStudy (raffle_snaffle):

Here is my work for 11a

OpenStudy (mathmale):

cumulative =?= continuous? No.

OpenStudy (raffle_snaffle):

cumulative and density? I have the formulas. I can do by hand. I don't know yet. I have not looked at solving these in my calculator. For parts C and D you can use a table in the back of the text. Lets move on to part b, c, and d

OpenStudy (mathmale):

First, I'd like to test y ou. If n=25, p=0.25 and x=3, what is p(x)?

OpenStudy (raffle_snaffle):

got PX=3) = 0.06410

OpenStudy (raffle_snaffle):

P(X=3)

OpenStudy (mathmale):

Looks OK, but I want to check; hold just a few sec, please.

OpenStudy (mathmale):

Good. You've got it.

OpenStudy (mathmale):

Just a sec, please

OpenStudy (mathmale):

I am skipping (a) and starting with (b) at your request, OK? (b) What is the probability that exactly 4 of the calls involve a fax message? Find binompdf(25,.25,4), please.

OpenStudy (mathmale):

This applies to "exactly 4."

OpenStudy (raffle_snaffle):

P(X=4) = 0.11756

OpenStudy (mathmale):

great. Part (b) done.

OpenStudy (mathmale):

Explain in your own words what "(c) What is the probability that at least 4 of the calls involve a fax message?" means. Could you give an example?

OpenStudy (raffle_snaffle):

P(X>=4)

OpenStudy (raffle_snaffle):

um... thinking

OpenStudy (raffle_snaffle):

are we looking above 4 so looking at 5,6,7,8,9,10,..... 25

OpenStudy (mathmale):

In this case, no. "at least 4" includes 4, as well as all the other integers up to and including 25.

OpenStudy (mathmale):

"at least 4" has a diff meaning than "above 4."

OpenStudy (raffle_snaffle):

most and least are the two I get confused with the most.

OpenStudy (raffle_snaffle):

least is implying we are looking at 4 and below?

OpenStudy (mathmale):

No, "4 or greater, up to and possibily including 25."

OpenStudy (mathmale):

I want you to buy me at least 1 ice cream cone of thanks, but preferably 2.

OpenStudy (mathmale):

0 ice cream cones would be a deal breaker. 1 cone would be OK (minimally acceptable) 2 cones would be great.

OpenStudy (raffle_snaffle):

this is implying X<=4 we are looking at random variable X above starting with 4.... so 4,5,6,7,8.... 25

OpenStudy (mathmale):

Exactly. Rather than calculate the probabilities for 4 through 25, we could calculate the probs P(x=0), P(x=1), etc., up thru P(x=3). That's 4 calculations. Or, y ou could find the cumulative probability for 0 thru 3 calls. Know how to do that on your TI-89?

OpenStudy (mathmale):

I get 0.09621. Most imp't is that you know how to use binompdf( ) and binomcdf( ).

OpenStudy (raffle_snaffle):

I said above... Lol you just implied below 0,1,2,3

OpenStudy (mathmale):

Nope, I've implied the prob. that you'd get 0, 1, 2 or 3 calls. I got 0.09621, which y ou should have gotten also. Now subtract this result from 1, please tell me your result, and please explain why we're subtracting here.

OpenStudy (raffle_snaffle):

P(X>=4) = 1 - P(X=0) + P(X=1) + P(X=2) + P(X=3)

OpenStudy (raffle_snaffle):

^^ I don't get this statement. We are subtracting the 1 because it's the complement. it represnts the 100%

OpenStudy (mathmale):

Actually, friend, that'd be 1-(P(X=0) + P(X=1) + P(X=2) + P(X=3))

OpenStudy (raffle_snaffle):

yes...

OpenStudy (mathmale):

Yes. you want the prob. of "at least 4." That covers most of the cases we're looking at, right? meaning that this prob is close to 1. We cheat a bit, start with 1, and then subtr. P(3 or fewer). Result is the prob. you wanted, that of 4 or more calls.

OpenStudy (raffle_snaffle):

can you draw this?

OpenStudy (mathmale):

1 => all cases .09 something => P(fewer than 4) (1-0.09 something) => P (4 or more calls are fax calls)

OpenStudy (mathmale):

I've cheated again by drawing this verbally.

OpenStudy (mathmale):